OpenStudy (anonymous):

The maximum value of |z1 - z2| satisfying the condition |z1| = 10 and |z2+3+4i| = 5 is?

OpenStudy (experimentx):

this is a circle centered at -3, 4 |z2+3+4i| = 5

OpenStudy (helder_edwin):

Let \(z_1=a+ib\) and \(z_2=c+id\). Then \[ \large 10=|z_1|=\sqrt{a^2+b^2}\Rightarrow100=a^2+b^2 \]

OpenStudy (experimentx):

|z1| = 10 <-- this is a circle centered at 0,0 with radius 10

OpenStudy (experimentx):

what's the max?? what's the max?

OpenStudy (anonymous):

Can we use triangle Inequality Here...on is there any easy method...

OpenStudy (experimentx):

* -3, -4

OpenStudy (anonymous):

It is the Diameter

OpenStudy (experimentx):

yep

OpenStudy (anonymous):

20....)

OpenStudy (anonymous):

yes 20